OpenStudy (anonymous):

f(x+4) shifts the function f(x), 4 units left on negative x axis. f(x)+4 shifts the function f(x), 4 units up on positive y axis. f(4x) will 1/4th the y coordinate only.( x remains as it is). 4f(x) will increase the y coordinate 4 times(x remains as it is).

OpenStudy (anonymous):

medal for anyone who just tells me

OpenStudy (anonymous):

i need this done lol i failed it once already

OpenStudy (anonymous):

@freckles

OpenStudy (anonymous):

@Mehek14

OpenStudy (anonymous):

@jabez177

OpenStudy (anonymous):

lol ill help

OpenStudy (anonymous):

would you be willing to just give me the answer or no?

OpenStudy (anonymous):

dont really now much on this but i will try my best can u put it in simpler terms and where are u from if u dont mind me asking

OpenStudy (anonymous):

Given the function f(x) = 3(x − 3)2 + 2, indicate the shifts that will affect the location of the vertex, and explain what effect they will have. Use complete sentences. f(x+4) f(x) + 4 f(4x) 4•f(x)

OpenStudy (anonymous):

i meant this lmfao

OpenStudy (anonymous):

lol wow whas about to say

OpenStudy (anonymous):

lol im sorry but can you please give me the answer to that or no

OpenStudy (anonymous):

yes i will here

Nnesha (nnesha):

`f(x−h)` here h is horizontal shift `f(x) + k` here k is vertical shift y=a(x-h)^2+k if h is negative graph will shift h unit to the right if h is positive graph will shift h unit to the left if k is positive graph shift k unit p if k is negative some unit down

OpenStudy (anonymous):

that's right?

Nnesha (nnesha):

what ?

OpenStudy (anonymous):

dang nnesha was gonna give her the answer

OpenStudy (anonymous):

Basically the graph of f(x) is the graph of the expression 6x^2 shifted 3 units to right of the origin and 2 units vertically from origin along the +ve direction of y axis. So vertex of the graph of f(x) will be at the point (3,2)

Nnesha (nnesha):

no i will not trust me.

OpenStudy (anonymous):

am i right

OpenStudy (anonymous):

hello